OpenStudy (anonymous):

find the constant variation k. m varies directly with n; m = 56 when n =8 k =?

OpenStudy (anonymous):

56/8=? OMG math!

Directrix (directrix):

7

OpenStudy (anonymous):

similarly, m = kn -> k = m/n = 56/8 = 7